Find the equation of the line which satisfies each set of conditions:
passes through the point with coordinates (-2, 1) and (0, 0)

Answer by venugopalramana(3286)

(-2, 1) and (0, 0)
The eqn.of a straight line through 2 points (x1,y1),(x2,y2)is given by
(y-y1)=[(y2-y1)/(x2-x1)]*(x-x1)..taking x1=0,y1=0 and x2=-2,y2=1..we get
y=(1/-2)*x or -2y=x or x+2y=0